Mario Brothers, a game manufacturer, has a new idea for an adventure game. It can either market the game as a traditional board game or as an interactive DVD, but not both. Consider the following cash flows of the two mutually exclusive projects. Assume the discount rate for both projects is 10 percent.

Year Board Game DVD

0 –$ 1,350 –$ 3,000

1 720 1,900

2 1,100 1,600

3 240 950

a. What is the payback period for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Payback period Board game = ____ years

Payback Period DVD=________ years

b. What is the NPV for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

NPV Board game= _______$

NPV DVD=________ $

c. What is the IRR for each project? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

IRR Board game=_______ %

IRR DVD=_______ %

d. What is the incremental IRR? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Incremental IRR=________%
Answer #1

a. Payback period is the time taken to recover the original investment of the project:

Payback for Board Game:

= 1 + $630/ $1100

= 1.5727

=1.57 ( rounded off to two decimal places)

Payback for DVD is :

= 1 + 1100/ 1600

= 1.6875

= 1.69 ( rounded off to two decimal places)

b. NPV of Board game :

($1350 ) + 720/1.1 + 1100/1.1^2 + 240/1.1^3

= $393.95

Similarly, for DVD the NPV is :

= $763.33

c.The IRR for Board game:

($1,350 ) + 720/ (1 + IRR)^1 + 1100/ (1 + IRR)^2 + 240/(1 + IRR)^3 = 0

So, the IRR is 27.9%

Similarly, the IRR for DVD is :

= 25.76%

d.Incremental IRR is :

CF0 = ($1650)

CF1= $1180

CF2= $500

CF3= $710

($1650 ) + 1180/(1 + IRR)^1 + 500/ (1 + IRR)^2 + 710/ (1 + IRR)^3 = 0

So, the incremental IRR is 23.96%